# ProSparse: Introducing and Enhancing Intrinsic Activation Sparsity within Large Language Models

### v4

@@ Abstract
Activation sparsity refers to the existence of considerable weakly-contributed elements among activation outputs. As a prevalent property of the models using the ReLU activation function, activation sparsity has been proven a promising paradigm to boost model inference efficiency. Nevertheless, most large language models (LLMs) adopt activation functions without intrinsic activation sparsity (e.g., GELU and Swish). Some recent efforts have explored introducing ReLU or its variants as the substitutive activation function to help LLMs achieve activation sparsity and inference acceleration, but few can simultaneously obtain high sparsity and comparable model performance. This paper introduces a simple and effective sparsification method named “ProSparse” to push LLMs for higher activation sparsity while maintaining comparable performance. Specifically, after substituting the activation function of LLMs with ReLU, ProSparse adopts progressive sparsity regularization with a factor smoothly increasing along the multi-stage sine curves. This can enhance activation sparsity and mitigate performance degradation by avoiding radical shifts in activation distributions. With ProSparse, we obtain high sparsity of 89.32% for LLaMA2-7B, 88.80% for LLaMA2-13B, and 87.89% for end-size MiniCPM-1B, respectively, achieving comparable performance to their original Swish-activated versions. These present the most sparsely activated models among open-source LLaMA versions and competitive end-size models, considerably surpassing ReluLLaMA-7B (66.98%) and ReluLLaMA-13B (71.56%). Our inference acceleration experiments further demonstrate the significant practical acceleration potential of LLMs with higher activation sparsity, obtaining up to 4.52$\times$ inference speedup.
@@ 1 Introduction
Recent years have witnessed the significant breakthrough made by large language models (LLMs), and these LLMs display commendable performance across a wide range of NLP tasks [8, 77, 54, 53, 72, 73, 1]. Nevertheless, the formidable computational costs required by the deployment and inference of LLMs pose a considerable challenge to the wider application of LLMs [3, 56]. The utilization of activation sparsity is one of the most promising techniques to reduce computational costs and enhance inference efficiency [44, 66], by discarding the redundant computation associated with the elements among LLM activation outputs that contribute weakly to the final result.
The adoption of ReLU, which naturally outputs zero elements, as the activation function is a straightforward method to achieve intrinsic activation sparsity and widely adopted in early LLMs [58, 86]. However, recent LLMs predominantly favor non-ReLU activation functions, such as GELU and Swish [72, 12, 2]. Although these non-ReLU LLMs may also display activation sparsity according to recent findings [88], such sparsity is manually imposed by searching adaptive activation thresholds (i.e., non-intrinsic), which can potentially lose minor neuron outputs and degrade performance. To pursue the sparsity-based inference acceleration, the task of ReLUfication is proposed [87, 50], aiming to introduce ReLU-based intrinsic activation sparsity into non-ReLU LLMs. Preliminary ReLUfication methods [87, 88] directly substitute the activation functions of non-ReLU LLMs with ReLU. As such substitution cannot overcome the inherent limitation imposed by the original non-ReLU activation distribution, inserted and shifted ReLU functions [50] are further introduced to enforce higher sparsity through radically shifting the activation distribution. Despite the promise of ReLUfication, existing efforts fail to achieve satisfactory activation sparsity and risk performance degradation caused by ReLUfication.
@@
In this paper, we propose a simple and effective ReLUfication method to help non-ReLU LLMs obtain high activation sparsity without performance degradation. We name the method “ProSparse”, which includes three steps shown in Figure 1: activation function substitution, progressive sparsity regularization, and activation threshold shifting. The first step is to replace the activation function of non-ReLU LLMs with ReLU and then apply continual training for adapting LLMs to the new ReLU activation. As discussed above, this can hardly achieve satisfactory sparsity. Therefore, in the second step, we apply sparsity regularization [47] to the intermediate activation outputs of the feed-forward networks (FFNs) within LLMs to seek higher activation sparsity. Considering the potential performance risks of forcing the fixed regularization factor [47, 41], we progressively increase the regularization factor in multiple stages. Concretely, the factor is set to a low constant value for the warmup stage. Next, during each subsequent stage, the factor undergoes a gradual increase along a gentle sine curve. Such progressive sparsity regularization can provide more time for the model to adapt to increasing regularization and avoid a radical shift in activation distribution, thereby mitigating performance degradation. The final step adopts FATReLU [38] to modify the vanilla ReLU by shifting its activation threshold to a positive value. This prunes less influential neurons to further improve sparsity.
@@
In experiments, we apply ProSparse to the ReLUfication of LLaMA2 [73] and end-size MiniCPM [34]. Activation sparsity of 89.32%, 88.80%, and 87.89% are successfully achieved for LLaMA2-7B, LLaMA2-13B, and MiniCPM-1B, respectively, with performance comparable to their original Swish-activated versions on various LLM benchmarks. Furthermore, we demonstrate the practical inference acceleration effect of higher activation sparsity, by respectively applying an approximate algorithm and an accurate algorithm to the inference of models with different sparsity. For the approximate one, we use PowerInfer [66], which achieves state-of-the-art speedup ratios tailored for sparsely activated LLMs at the expense of potentially inaccurate inference due to the mistakes of activation predictors. For the accurate one, we implement and release two GPU operators that leverage the input-side and output-side sparsity during the computation of ReLU-activated FFN layers [^1] Source codes for these two operators are available at https://github.com/Raincleared-Song/sparse_gpu_operator..
@@
In summary, we make the following contributions in this paper: (1) We propose ProSparse, an effective ReLUfication method that converts non-ReLU LLMs into much sparser ReLU-activated LLMs without performance degradation. (2) Sparsely activated versions of LLaMA2-7B, LLaMA2-13B, and MiniCPM-1B comparable to their original Swish-activated versions in performance are both obtained and available [^2] Models are available at https://huggingface.co/SparseLLM/prosparse-llama-2-7b, https://huggingface.co/SparseLLM/prosparse-llama-2-13b, and https://huggingface.co/openbmb/MiniCPM-S-1B-sft, respectively.. (3) We demonstrate the practical inference acceleration effect of higher activation sparsity that ProSparse can reach. Valuable observations and analyses are also conducted.
@@ 2 Preliminaries and Related Works
Here we mainly introduce works on improving LLM inference efficiency. Refer to existing surveys [7, 91] for more details about LLMs. Related works about $L_{1}$ regularization are listed in Appendix A.
@@
Despite the commendable performance of LLMs, the sustainable increase in model scales brings the exponential growth of inference computations, making the deployment of LLMs a formidable challenge [37, 44]. To reduce the computational costs required by LLM inference, various model compression methods have been proposed, such as quantization [25, 35, 52, 90, 5, 82, 83], pruning [25, 26, 51, 31, 48, 67, 21, 81], and distillation [30, 69, 71, 24, 33]. Efficient sampling methods have also been proposed to achieve faster inference decoding speed [39, 75, 9, 49]. In general, none of these acceleration methods leverages the intrinsic mechanisms within LLMs.
@@
Recent works [43, 44, 66] have noticed the intrinsic activation sparsity within some LLMs and its potential in inference acceleration. Activation sparsity refers to the phenomenon where considerable zero or negligible elements in activation outputs, corresponding to certain model parameters (i.e., neurons), have a weak impact on LLM outputs given a specific input. These weakly-contributed parameters are regarded as inactivated and can thus be skipped during inference to save considerable computational resources. Notably, the utilization of activation sparsity is orthogonal to model compression and efficient sampling, and these approaches can be readily accumulated.
Difference between Activation Sparsity and Pruning
Generally, pruning realizes sparsity by removing certain elements (e.g., neurons, weights, or structured blocks) in LLMs. However, the sparsity introduced by pruning is statically limited to model weights and independent of the inputs. High static sparsity is often accompanied by considerable performance degradation compared to the original dense model [21, 81]. By contrast, activation sparsity is dynamically determined by the input data and thus potentially compromises less model capacity and downstream task performance.
@@
Previous studies have marked the existence of activation sparsity in almost any ReLU-activated Transformer architecture [43], from LLMs (e.g., T5 [58] and OPT [86]) to vision models (e.g., ViT [19]). However, recent LLMs such as Falcon [2] and LLaMA [73] prevalently adopt non-ReLU activation functions such as GELU [29] and Swish [20] without intrinsic activation sparsity. Therefore, to leverage the merits of activation sparsity without training a ReLU-activated LLM from scratch, many works conduct ReLUfication, the task of introducing sparse ReLU-based activations into non-ReLU LLMs. In MoEfication [87], a GELU-activated BERT [17] is converted into a ReLU-activated version through direct activation function substitution and additional training. ReluLLaMA and ReluFalcon apply a similar paradigm to Falcon and LLaMA, respectively [88]. Since activation substitution cannot reach satisfactory sparsity, mainly due to the unhandled limitation of the original non-ReLU activation distribution, the inserted and shifted ReLU activation functions are introduced [50], conducting a radical shift in activation distribution. Although these operations are claimed to achieve sparsity of nearly 95%, we cannot replicate the results in our experiments (see the 3rd paragraph of Section 4.4) and the sparsity is still limited.
As discussed above, we can clearly recognize the promise of activation sparsity and also observe the key challenge of leveraging ReLUfication to achieve activation sparsity: how to achieve high sparsity and mitigate performance degradation concurrently. These are what ProSparse is designed for.
@@ 3.1 Definitions and Notations
For the convenience of subsequent demonstrations, here we define activation sparsity within LLMs in detail. Since the activation function mainly exists in the FFN layers of LLMs, we first formalize the computation process of FFNs. Given the hidden dimension $d_{model}$ and the FFN intermediate dimension $d_{ff}$, the computation process of a gated FFN (i.e., the most widely adopted FFN architecture in recent Transformer-based LLMs [16, 64]) can be formalized as:
$\mathbf{s}=\sigma(\mathbf{x}\mathbf{W}_{s}^{T}),\quad\mathbf{x}_{1}=\mathbf{s}\odot(\mathbf{x}\mathbf{W}_{1}^{T}),\quad\text{FFN}(\mathbf{x})=\mathbf{x}_{1}\mathbf{W}_{2}^{T},$ (1)
@@
We define the activation sparsity (hereinafter abbreviated as sparsity) as the ratio of zero elements (i.e., inactivated elements) in $\mathbf{x}_{1}$ for a specific input $\mathbf{x}$. Since the sparsity varies in different layers for different inputs, we evaluate the sparsity of an LLM using the average sparsity, defined as the average value of sparsity across all layers in an LLM on a sufficiently large amount of input data.
In this paper, we focus on the task of ReLUfication, namely converting an LLM using a non-ReLU activation function $\sigma$ (e.g., GELU or Swish) into a ReLU-activated one, while making the activation sparsity as high as possible and mitigating performance degradation.
@@ 3.2 ProSparse
We propose ProSparse to achieve the above targets. In ProSparse, three steps are carefully designed to introduce and enhance the intrinsic activation sparsity for a non-ReLU LLM: (1) activation function substitution; (2) progressive sparsity regularization; (3) activation threshold shifting.
@@
Nevertheless, activation function substitution by nature does not change the activation distribution, which will potentially limit the sparsity to relatively low values. To push for higher sparsity, a typical method is $L_{1}$ sparsity regularization [43], which introduces the $L_{1}$ regularization loss as an extra training target. Given the intermediate output $\mathbf{x}_{1}$ of the $i$-th FFN layer in an LLM, the regularization loss is defined as:
@@
Considering the potentially unstable training and performance degradation due to fixed regularization factors [23, 38, 43], we propose the progressive sparsity regularization, where the factor $\lambda$ is carefully scheduled to gently increase in multiple stages, as displayed in Algorithm 3.2.

### v5

@@ Abstract
Activation sparsity refers to the existence of considerable weakly-contributed elements among activation outputs, serving as a promising paradigm for accelerating model inference. Nevertheless, most large language models (LLMs) adopt activation functions without intrinsic activation sparsity (e.g., GELU and Swish). Some recent efforts have explored introducing ReLU or its variants as the substitutive activation function to pursue activation sparsity and acceleration, but few can simultaneously obtain high activation sparsity and comparable model performance. This paper introduces a simple and effective method named “ProSparse” to sparsify LLMs while achieving both targets. Specifically, after introducing ReLU activation, ProSparse adopts progressive sparsity regularization with a factor smoothly increasing for multiple stages. This can enhance activation sparsity and mitigate performance degradation by avoiding radical shifts in activation distributions. With ProSparse, we obtain high sparsity of 89.32% for LLaMA2-7B, 88.80% for LLaMA2-13B, and 87.89% for end-size MiniCPM-1B, respectively, with comparable performance to their original Swish-activated versions. These present the most sparsely activated models among open-source LLaMA versions and competitive end-size models. Inference acceleration experiments further demonstrate the significant practical acceleration potential of LLMs with higher activation sparsity, obtaining up to 4.52$\times$ inference speedup.
ProSparse: Introducing and Enhancing Intrinsic Activation Sparsity within Large Language Models
Chenyang Song1, Xu Han1∗, Zhengyan Zhang1, Shengding Hu1, Xiyu Shi2 Kuai Li3, Chen Chen3, Zhiyuan Liu1thanks: Corresponding author: Xu Han (han-xu@tsinghua.edu.cn) and Zhiyuan Liu (liuzy@tsinghua.edu.cn) , Guangli Li2, Tao Yang3, Maosong Sun1 1Dept. of Comp. Sci. & Tech., Institute for AI, Tsinghua University, Beijing, China 2SKLP, Institute of Computing Technology, Chinese Academy of Sciences, China 3Tencent Machine Learning Platform, China scy22@mails.tsinghua.edu.cn, han-xu@tsinghua.edu.cn
@@ 1 Introduction
Recent years have witnessed significant breakthroughs made by large language models (LLMs) with commendable performance across a wide range of NLP tasks Brown et al. (2020); Wei et al. (2021); Ouyang et al. (2022); OpenAI (2023); Achiam et al. (2023). Nevertheless, the formidable computational costs required by the deployment and inference of LLMs pose a considerable challenge Aminabadi et al. (2022); Pope et al. (2023). The utilization of activation sparsity is one of the most promising techniques to enhance inference efficiency Liu et al. (2023); Song et al. (2023), by discarding the redundant computation associated with the elements among LLM activation outputs that contribute weakly to the final outputs.
@@
The adoption of ReLU, which naturally outputs zero elements, as the activation function is a straightforward method to achieve intrinsic activation sparsity in early LLMs Raffel et al. (2020); Zhang et al. (2022a). However, recent LLMs predominantly favor non-ReLU activation functions, such as GELU and Swish Touvron et al. (2023a); Chowdhery et al. (2023); Almazrouei et al. (2023). Although these non-ReLU LLMs may also display activation sparsity Zhang et al. (2024), such sparsity is manually imposed by searching adaptive activation thresholds (i.e., non-intrinsic), which can potentially lose minor neuron outputs and degrade performance. To pursue the sparsity-based inference acceleration, the task of ReLUfication is proposed, aiming to introduce ReLU-based intrinsic activation sparsity into non-ReLU LLMs. Preliminary methods Zhang et al. (2022b, 2024) directly substitute the activation functions with ReLU. As such substitution cannot overcome the inherent limitation imposed by the original dense activation distribution, inserted and shifted ReLU functions Mirzadeh et al. (2023) are introduced to enforce higher sparsity through radically shifting the activation distribution. However, existing efforts fail to achieve satisfactory sparsity and risk performance degradation.
@@
In this paper, we propose a simple and effective ReLUfication method named “ProSparse” to help non-ReLU LLMs obtain high activation sparsity without performance degradation. ProSparse includes three steps shown in Figure 1: activation function substitution, progressive sparsity regularization, and activation threshold shifting. The first step is to replace the activation function with ReLU and then apply continual training. As discussed above, this can hardly achieve satisfactory sparsity. Therefore, in the second step, we apply sparsity regularization Ma et al. (2019) to the intermediate activation outputs of the feed-forward networks (FFNs) within LLMs to seek higher sparsity. Considering the potential performance risks of forcing the fixed regularization factor Ma et al. (2019); Li et al. (2020), we progressively increase the factor in multiple stages, including one flat warmup stage and multiple incremental stages along gentle sine curves. Such progressive regularization can provide more time for adaption to increasing regularization and avoid a radical shift in activation distribution, thereby mitigating performance degradation. The final step adopts FATReLU Kurtz et al. (2020), shifting the ReLU activation threshold to a positive value. This prunes less influential neurons to improve sparsity.
@@
In experiments, we apply ProSparse to the ReLUfication of LLaMA2 Touvron et al. (2023b) and end-size MiniCPM Hu et al. (2024). Activation sparsity of 89.32%, 88.80%, and 87.89% are successfully achieved for LLaMA2-7B, LLaMA2-13B, and MiniCPM-1B, respectively, with performance comparable to their original Swish-activated versions on various LLM benchmarks. Furthermore, we demonstrate the practical inference acceleration of higher activation sparsity, by respectively applying an approximate algorithm and an accurate algorithm to the inference of models with different sparsity. For the approximate one, we use PowerInfer Song et al. (2023), which achieves state-of-the-art speedup ratios tailored for sparsely activated LLMs but risks inaccurate inference due to the mistakes of activation predictors. For the accurate one, we implement and release two GPU operators that leverage the input-side and output-side sparsity during the computation of ReLU-activated FFNs [^1] Source codes for these two operators are available at https://github.com/Raincleared-Song/sparse_gpu_operator..
@@
In summary, we make the following contributions in this paper: (1) We propose ProSparse, an effective ReLUfication method that converts non-ReLU LLMs into much sparser ReLU-activated LLMs without performance degradation. (2) Sparsely activated versions of LLaMA2-7B, LLaMA2-13B, and MiniCPM-1B comparable to their original Swish-activated versions in performance are both obtained and available [^2] Models are respectively available at https://huggingface.co/SparseLLM/prosparse-llama-2-7b, https://huggingface.co/SparseLLM/prosparse-llama-2-13b, and https://huggingface.co/SparseLLM/ProSparse-MiniCPM-1B-sft.. (3) We demonstrate the practical inference acceleration effect of higher activation sparsity that ProSparse can reach. Valuable observations and analyses are also conducted.
@@ 2 Preliminaries and Related Works
Here we discuss how to improve LLM inference efficiency. Refer to existing surveys Zhao et al. (2023) for works about LLMs and Appendix A for works about $L_{1}$ regularization.
@@
Efficiency has long been a crucial topic in various AI applications Chen et al. (2023b). The sustainable increase in LLM scales brings the exponential growth of inference computations, making the deployment of LLMs a formidable challenge Kaplan et al. (2020); Liu et al. (2023). To reduce the computational costs required by LLM inference, various model compression or decoding acceleration methods have been proposed, such as quantization Jacob et al. (2018); Nagel et al. (2019); Zhao et al. (2019); Bai et al. (2022); Xiao et al. (2023); Yao et al. (2023), pruning Hoefler et al. (2021); Ma et al. (2023); Sun et al. (2023); Frantar and Alistarh (2023); Xia et al. (2023), distillation Tang et al. (2019); Touvron et al. (2021); Gu et al. (2023); Hsieh et al. (2023), and efficient sampling methods Leviathan et al. (2023); Wang et al. (2023); Chen et al. (2023a); Miao et al. (2023). While these works have proved effective for inference acceleration and other scenarios (e.g., secure federated learning Ding et al. (2023b)), none of these methods leverages the intrinsic mechanisms within LLMs.
@@
Recent works Li et al. (2022); Liu et al. (2023); Song et al. (2023) have noticed the intrinsic activation sparsity within some LLMs and its potential in inference acceleration. Activation sparsity refers to the phenomenon where considerable zero or negligible elements in activation outputs, corresponding to certain model parameters (i.e., neurons), have a weak impact on LLM outputs given a specific input. These weakly-contributed parameters are regarded as inactivated and can thus be skipped during inference to save computational resources. Notably, the utilization of activation sparsity is orthogonal to model compression and efficient sampling, and these approaches can be readily combined. Another fact worth attention is the fundamental difference between activation sparsity and pruning, see Appendix A.
@@
Activation sparsity naturally exists in ReLU-activated architecture Li et al. (2022), from LLMs Raffel et al. (2020); Zhang et al. (2022a) to vision models Dosovitskiy et al. (2020). However, recent LLMs such as Falcon Almazrouei et al. (2023) and LLaMA Touvron et al. (2023b) prevalently adopt non-ReLU activation functions such as GELU Hendrycks and Gimpel (2016) and Swish Elfwing et al. (2018) without intrinsic activation sparsity. Therefore, to leverage the merits of activation sparsity without training a ReLU-activated LLM from scratch, many works conduct ReLUfication, introducing sparse ReLU-based activations into non-ReLU LLMs. Zhang et al. (2022b) converts a GELU-activated BERT Devlin et al. (2018) into a ReLU-activated version through activation function substitution and additional training. ReluLLaMA and ReluFalcon apply a similar paradigm to Falcon and LLaMA, respectively Zhang et al. (2024). Since activation substitution cannot reach satisfactory sparsity, mainly due to the unhandled limitation of the original dense activation distribution, the inserted and shifted ReLU activation functions are introduced Mirzadeh et al. (2023), conducting a radical shift in activation distribution. Although these operations are claimed to achieve sparsity of nearly 95%, we cannot replicate the results in our experiments (see the 3rd paragraph of Section 4.4) and the sparsity is still limited. By contrast, ProSparse is a ReLUfication method designed to achieve high sparsity and mitigate performance degradation concurrently.
@@ 3.1 Definitions and Notations
For the convenience of subsequent demonstrations, here we define activation sparsity in detail. Since the activation function mainly exists in the FFNs within LLMs, we first discuss the computation process of FFNs. Given the hidden dimension $d_{model}$ and the intermediate dimension $d_{ff}$, the computation process of a gated FFN (i.e., the most widely adopted FFN architecture in recent LLMs Dauphin et al. (2017); Shazeer (2020)) can be formalized as:
$\mathbf{s}=\sigma(\mathbf{x}\mathbf{W}_{s}^{T}),$ $\mathbf{x}_{1}=\mathbf{s}\odot(\mathbf{x}\mathbf{W}_{1}^{T}),$ (1)
$\text{FFN}(\mathbf{x})$ $=\mathbf{x}_{1}\mathbf{W}_{2}^{T},$
@@
We define the activation sparsity (hereinafter abbreviated as sparsity) as the ratio of zero elements (i.e., inactivated elements) in $\mathbf{x}_{1}$ for a specific input $\mathbf{x}$. The sparsity of an LLM is evaluated using the average sparsity, defined as the average value of sparsity across all layers in an LLM on a sufficiently large amount of input data.
In this paper, we focus on the task of ReLUfication, namely converting an LLM using a non-ReLU activation function $\sigma$ (e.g., GELU or Swish) into a ReLU-activated one, while making the average sparsity as high as possible and mitigating performance degradation.
@@ 3.2 ProSparse
We propose ProSparse to achieve the above targets. Three steps are carefully designed to introduce and enhance the intrinsic activation sparsity for a non-ReLU LLM: (1) activation function substitution; (2) progressive sparsity regularization; (3) activation threshold shifting.
@@
Nevertheless, activation function substitution by nature does not change the activation distribution, which will potentially limit the sparsity to relatively low values. To push for higher sparsity, a typical method is $L_{1}$ sparsity regularization Li et al. (2022), which introduces the $L_{1}$ regularization loss as an extra training target. Given the intermediate output $\mathbf{x}_{1}$ of the $i$-th FFN layer in an LLM, the regularization loss is defined as:
@@
Considering the potential performance degradation due to fixed regularization factors Georgiadis (2019); Kurtz et al. (2020); Li et al. (2022), we propose the progressive sparsity regularization, where the factor $\lambda$ is carefully scheduled to gently increase in multiple stages. Refer to Appendix B for more details.